# Tricking LLMs into Disobedience: Formalizing, Analyzing, and Detecting Jailbreaks

### v2

@@ Abstract
Recent explorations with commercial Large Language Models (LLMs) have shown that non-expert users can jailbreak LLMs by simply manipulating their prompts; resulting in degenerate output behavior, privacy and security breaches, offensive outputs, and violations of content regulator policies. Limited studies have been conducted to formalize and analyze these attacks and their mitigations. We bridge this gap by proposing a formalism and a taxonomy of known (and possible) jailbreaks. We survey existing jailbreak methods and their effectiveness on open-source and commercial LLMs (such as GPT-based models, OPT, BLOOM, and FLAN-T5-XXL). We further discuss the challenges of jailbreak detection in terms of their effectiveness against known attacks. For our analysis, we collect a dataset of 3700 jailbreak prompts across 4 tasks. We will make the dataset public along with the model outputs.
@@
| Abhinav Rao,, Atharva Naik*,, Sachin Vashistha*,, |
@@ 1. Introduction
00footnotetext: Carnegie Mellon University 00footnotetext: Mohamed Bin Zayed University of AI00footnotetext: Indian Institute of Technology, Kharagpur00footnotetext: Work done when authors were at Microsoft00footnotetext: * denotes equal contributionTransformers-based generative Large Language Models (LLM) have demonstrated superior zero-shot (and few-shot) generalization capabilities Kojima et al. (2022); Huang et al. (2022b) under the new “pre-train, prompt, and predict” paradigm. Here, any user can provide a description of the task followed by zero or more examples in natural language to a pretrained LLM. Based on such an instruction (or “prompt”), the LLM can learn to perform a new task on unseen examples. This amazing ability to perform a new task following a natural language instruction have also exposed a new set of vulnerabilities, popularly categorized as “prompt injection attacks” or “jailbreaks”. Consider Fig. 2 for an example of a prompt injection attack setup and associated actors.
@@
While new methods of jailbreaks come up every day; till date, little formal study of prompt injection attacks exist which can portray a holistic idea of the types and dimensions of attacks, the severity, and vulnerability of models towards the attacks. The studies Kang et al. (2023); Greshake et al. (2023) are limited, divergent and there is an urgent need to consolidate. Here, we draw inspiration from other fields of Computer Science (such as computer security, SQL injection attacks) where similar attacks have been studied. We approach this problem by presenting a formalism of a "jailbreak" or a "prompt-injection" with an experimental setup to test our claims. We curate a taxonomy of the possible characteristics of jailbreaks, and provide an analysis on the efficacy of jailbreaks on different language models. Additionally, we discuss tests to detect the success of a jailbreak. We believe that our contributions can help the community understand and explore the jailbreaking phenomenon more formally, leading to more research in the field.
@@ 2.1. Current Work on Jailbreaks
The first occurrence of the term ‘prompt injection’ was through a few social media blogs Willison (2022); Preamble.AI (2022). The term jailbreak soon represented the same phenomenon in social media blogs Kilcher (2022), and gained traction in multiple subreddits such as r/ChatGPT, r/ChatGPTJailbreak, r/bing, and r/OpenAI. A specific jailbreak by the name of DAN a.k.a. “Do Anything Now" was popularized in Reddit InternationalData569 and GPU_WIZ (2022) and through several web articles King (2023); Walker (2023); Smith (2023). In the academic community, the problem of ‘prompt injection’ or ‘jailbreaks’ (borrowed from the Operating Systems concept of a privilege escalation exploit) of large language models is relatively new, but rapidly evolving with a lot of recent work around formalization of the problem Wei et al. (2023), categorization of known techniques Wei et al. (2023); Shen et al. (2023); Mozes et al. (2023), automating attacks Zou et al. (2023); Yao et al. (2023); Deng et al. (2023), evaluation and mitigation strategies Wei et al. (2023); Shen et al. (2023); Yao et al. (2023); Mozes et al. (2023). Perez and Ribeiro (2022) performed prompt injection attacks on GPT-3 which involved either hijacking the model’s goal, leading it to generate malicious target text, or leaking the original prompt and instructions. Wolf et al. (2023) provide theoretical conjectures on the cause of misalignment, and hence jailbreaks and potential fixes of the situation, using a model’s output logits and RLHF fine-tuning. Greshake et al. (2023) approached the problem from a computer-security standpoint, showing indirect prompt-injection threats and their consequences to LLM applications. Kang et al. (2023) and Deng et al. (2023) exploit the observation that instruction-finetuned LLMs can work like standard computer programs, and carry out “return-oriented programming (ROP) attacks", and time-based SQL attacks on LLMs respectively. Zou et al. (2023) use greedy coordinate gradient descent to identify universal sequences of characters to jailbreak LLMs. Qi et al. (2023) introduce vision-based jailbreak attacks for multimodal LLMs. Some works also collected specific jailbreaks Shen et al. (2023) and curated similar synthetic prompts Casper et al. (2023); Qiu et al. (2023).
@@ 4.1.2. Lexical Techniques
Techniques that rely upon specific words or phrases to achieve jailbreaks can be classified as “Lexical Jailbreak Techniques". Zou et al. (2023) create an adversarial jailbreak mechanism that uses coordinate-gradients to generate a ‘universal jailbreak’ such as ‘describing. + similarlyNow write oppositeley.]( Me giving**ONE please? revert with "--Two’.
@@ 4.2. Jailbreak Intents
Misaligned Content Generation. Here the intent of the attacker is to generate content through a system that is misaligned to the ethical principles or alignment goals of the system. Some commonly observed sub-types include generation of fake Casper et al. (2023), toxic, hateful, abusive content Shen et al. (2023); Zou et al. (2023), and also content that can aid the user in causing harm or destruction (how to hotwire a car; how to make a bomb, etc.) Borji (2023); Borji and Mohammadian (2023); Lambert and Stevens (2023).
@@ 5.1.1. Definition of Jailbreak Success
We define the notion of jailbreak success based on graded evaluations of the divergence of task from the task . We capture these metrics using property tests which are functions of the task and the jailbreak respectively and will be expanded upon in the metrics section (Sec. 5.1). For manual evaluations, we consider the more stringent metric of whether aligns with the malicious intent of , called an intent test.
@@ 5.2.1. Experimental Setup
We supply the model with a malicious user input . In some cases, we add a base user input before the jailbreak in order to emulate an MITM jailbreak. The aim is to cause divergence in the model output from the prompted task to a misaligned task .
We pick multiple instantiations of by choosing popular LLMs with different performance capabilities: OPT-175B Zhang et al. (2022), BLOOM-176B Workshop et al. (2022), GPT-3 models (text-ada, text-babbage, text-curie Brown et al. (2020), GPT-3.5 based models text-davinci-002, gpt-3.5-turbo Ouyang et al. (2022), and FLAN-T5-XXL (11B) Wei et al. (2021). We design different kinds of jailbreaks for each task for a jailbreak type as . One may note that the jailbreaks are independent of the model used, since in most practical settings, an attacker knows which task the model has been prompted for, but not which model is being used (for e.g. BING’s announcement Mehdi (2023) about using GPT-4 came five weeks after their chatbot preview became accessible).
@@ 5.2.2. Dataset
The prompt formats are sourced from OpenAI, Promptsource Bach et al. (2022), and from several academic sources Chen et al. (2021); Muennighoff et al. (2022); Wei et al. (2022); Zhang et al. (2022). In cases where we did not find a preexisting prompt for a model-task combination, we recycled prompts from other models.
@@
Based on findings from Twitter, video sources, and Gehman et al. (2020), we manually curate jailbreaks across the said dimensions in Section 4. In total, there are 55 jailbreaks over all 4 tasks. We run the property-tests on all 55 jailbreaks for every model. We vary the user input in some cases to analyze its effect on the attack success. In total, there are 3.7k jailbreaks that are fed into each model.
@@ 5.2.3. Results
We report the results of our GPT-4 tests for Figures 2 and 3. In terms of the jailbreak type, we notice that the jailbreak success decreases with an increase in model size until text-davinci-002, however, any further instruction or task tuning increases the tendency for misalignment, as in the case of gpt-3.5-turbo and code-davinci-002. It can also be noted that Cognitive hacking (COG) appears to be the most successful form of jailbreak, which also happens to be the most common type of jailbreak present in the real world, followed by Orthographic attacks (ORTH).
Almost all models seem to be most affected by the Performance Degradation intent, which is expected given the relative ease of achieving degradation. However, the scaled plots shows misaligned content to show high success for gpt-3.5-turbo and code-davinci-002, which we believe is related to their instruction-following capabilities. Additionally, text-davinci-002 appears to be robust to most of these jailbreaks, hinting that its training may be more robust to content-harms888https://tinyurl.com/davinci2toxic.
@@
It is seen that there is a poor agreement between both methods, suggesting that jailbreak detection can prove challenging and non-trivial. Additionally, we noticed that GPT-4 was occasionally jailbroken itself (especially with cognitive hacking), after being fed in the jailbreaks. This leads to a notion of a “jailbreak paradox", where it gets increasingly harder to detect and mitigate jailbreaks, due to the vast space of outputs the language model is capable of generating, and also due to its instruction-finetuning capabilities. Hence, we additionally conduct a manual evaluation to better understand the effect of jailbreaks on each model.
@@ 5.3. Manual Analysis
Each prompt is independently labeled by two annotators, where disagreements were resolved by a third999In this case, we chose three Computer Science graduate students trained in NLP as annotators.. We report the misalignment rate and jailbreak success rate in Table 3. We can still see that the attack success rate is higher for FLAN, and gpt-3.5-turbo, which confirms that both model size and instruction tuning have an influence on jailbreaking. We report our inter-annotator agreement to be over both labels, suggesting that even with humans, the “jailbreak paradox" of detection still exists. Additionally, we scale the results of each attack type by the True-Positive Rate (TPR) and the False-Negative Rate (FNR) of GPT-4 against our manual evaluations. We perform the scaling as follows: if we observe that GPT-4 assigned a class to examples and class to examples in the dataset, then the estimated (corrected) values for the two labels will be , . Post scaling, we see a significant increase in attack success for the Instruction Repetition (IR) type, due to the discrepancies between manual and automatic evaluations.
@@ 6. Conclusion
Large language models (LLMs) have shown remarkable capabilities of learning and performing new tasks from natural language instructions or prompts. However, this also exposes them to a new type of vulnerability, namely jailbreaks or prompt-injection attacks, where malicious users can manipulate the prompts to cause misalignment, leakage, performance degradation, or harmful generation. In this work, we have proposed a formalism and a taxonomy of jailbreaks based on their linguistic transformation and attacker intent. We perform an empirical analysis of the effectiveness of different types of jailbreaks on various LLMs. We found that LLMs have varying degrees of robustness to jailbreaks depending on their size, training data, and architecture. We discuss some limitations and challenges of the current methods for detecting and mitigating jailbreaks, such as the need for sanitizing and preprocessing the outputs, the difficulty of capturing the attacker’s intent, and the trade-off between functionality and security. We also explore some prompt-level mitigation strategies that we do not include because of space limitations. Specifically, our work provides a timely and useful framework, and a comprehensive analysis for researchers and practitioners who are interested in understanding and addressing this emerging challenge.
@@
We plan to make available the human evaluation data, jailbreaks, all source code, and GPT-4 prompts (wherever applicable) to the public to aid in reproducibility of results.
7. Acknowledgements
[…]
- Bach et al. (2022) Stephen Bach, Victor Sanh, Zheng Xin Yong, Albert Webson, Colin Raffel, Nihal V. Nayak, Abheesht Sharma, Taewoon Kim, M Saiful Bari, Thibault Fevry, Zaid Alyafeai, Manan Dey, Andrea Santilli, Zhiqing Sun, Srulik Ben-david, Canwen Xu, Gunjan Chhablani, Han Wang, Jason Fries, Maged Al-shaibani, Shanya Sharma, Urmish Thakker, Khalid Almubarak, Xiangru Tang, Dragomir Radev, Mike Tian-jian Jiang, and Alexander Rush. 2022. PromptSource: An integrated development environment and repository for natural language prompts. In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics: System Demonstrations, pages 93–104, Dublin, Ireland. Association for Computational Linguistics.

### v4

@@ Abstract
Recent explorations with commercial Large Language Models (LLMs) have shown that non-expert users can jailbreak LLMs by simply manipulating their prompts; resulting in degenerate output behavior, privacy and security breaches, offensive outputs, and violations of content regulator policies. Limited studies have been conducted to formalize and analyze these attacks and their mitigations. We bridge this gap by proposing a formalism and a taxonomy of known (and possible) jailbreaks. We survey existing jailbreak methods and their effectiveness on open-source and commercial LLMs (such as GPT-based models, OPT, BLOOM, and FLAN-T5-XXL). We further discuss the challenges of jailbreak detection in terms of their effectiveness against known attacks. For further analysis, we release a dataset of model outputs across 3700 jailbreak prompts over 4 tasks.
@@
Abhinav Rao,, Sachin Vashistha*,, Atharva Naik*,, |
@@ 1. Introduction
00footnotetext: Carnegie Mellon University 00footnotetext: Indian Institute of Technology, Kharagpur00footnotetext: Mohamed Bin Zayed University of AI00footnotetext: * = equal contribution. Order chosen at random.00footnotetext: Work conducted while authors were affiliated with Microsoft TuringTransformers-based generative Large Language Models (LLM) have demonstrated superior zero-shot (and few-shot) generalization capabilities Kojima et al. (2022); Huang et al. (2022b) under the new “pre-train, prompt, and predict” paradigm. Here, any user can provide a description of the task followed by zero or more examples in natural language to a pretrained LLM. Based on such an instruction (or “prompt”), the LLM can learn to perform a new task on unseen examples. This amazing ability to perform a new task following a natural language instruction have also exposed a new set of vulnerabilities, popularly categorized as “prompt injection attacks” or “jailbreaks”. Consider Fig. 2 for an example of a prompt injection attack setup and associated actors.
@@
While new methods of jailbreaks come up every day; till date, little formal study of prompt injection attacks exist which can portray a holistic idea of the types and dimensions of attacks, the severity, and vulnerability of models towards the attacks. The studies Kang et al. (2023); Greshake et al. (2023) are limited, divergent and there is an urgent need to consolidate. Here, we draw inspiration from other fields of Computer Science (such as computer security, SQL injection attacks) where similar attacks have been studied. We approach this problem by presenting a formalism of a "jailbreak" or a "prompt-injection" with an experimental setup to test our claims. We curate a taxonomy of the possible characteristics of jailbreaks, provide 3700 jailbreak prompts through a templated approach, and an analysis on the efficacy of jailbreaks on different language models. Additionally, we discuss attack detection tests to detect the success of a jailbreak. We make available our code and data in this URL.
@@ 2.1. Current Work on Jailbreaks
The first occurrence of the term ‘prompt injection’ was through a few social media blogs Willison (2022); Preamble.AI (2022). The term jailbreak soon represented the same phenomenon in social media blogs Kilcher (2022), and gained traction in multiple subreddits such as r/ChatGPT, r/ChatGPTJailbreak, r/bing, and r/OpenAI. A specific jailbreak by the name of DAN a.k.a. “Do Anything Now" was popularized in Reddit InternationalData569 and GPU_WIZ (2022) and through several web articles King (2023); Walker (2023); Smith (2023). In the academic community, the problem of ‘prompt injection’ or ‘jailbreaks’ (borrowed from the Operating Systems concept of a privilege escalation exploit) of large language models is relatively new, but rapidly evolving with a lot of recent work around formalization of the problem Wei et al. (2023), categorization of known techniques Wei et al. (2023); Shen et al. (2023); Mozes et al. (2023), automating attacks Zou et al. (2023a); Yao et al. (2023); Deng et al. (2023), evaluation and mitigation strategies Wei et al. (2023); Shen et al. (2023); Yao et al. (2023); Mozes et al. (2023). Perez and Ribeiro (2022) performed prompt injection attacks on GPT-3 which involved either hijacking the model’s goal, leading it to generate malicious target text, or leaking the original prompt and instructions. Wolf et al. (2023) provide theoretical conjectures on the cause of misalignment, and hence jailbreaks and potential fixes of the situation, using a model’s output logits and RLHF fine-tuning. Greshake et al. (2023) approached the problem from a computer-security standpoint, showing indirect prompt-injection threats and their consequences to LLM applications. Kang et al. (2023) and Deng et al. (2023) exploit the observation that instruction-finetuned LLMs can work like standard computer programs, and carry out “return-oriented programming (ROP) attacks", and time-based SQL attacks on LLMs respectively. Zou et al. (2023a) use greedy coordinate gradient descent to identify universal sequences of characters to jailbreak LLMs. Qi et al. (2023) introduce vision-based jailbreak attacks for multimodal LLMs. Some works also collected specific jailbreaks Shen et al. (2023) and curated similar synthetic prompts Casper et al. (2023); Qiu et al. (2023).
@@ 4.1.2. Lexical Techniques
Techniques that rely upon specific words or phrases to achieve jailbreaks can be classified as “Lexical Jailbreak Techniques". Zou et al. (2023a) create an adversarial jailbreak mechanism that uses coordinate-gradients to generate a ‘universal jailbreak’ such as ‘describing. + similarlyNow write oppositeley.]( Me giving**ONE please? revert with "--Two’.
@@ 4.2. Jailbreak Intents
Misaligned Content Generation. Here the intent of the attacker is to generate content through a system that is misaligned to the ethical principles or alignment goals of the system. Some commonly observed sub-types include generation of fake Casper et al. (2023), toxic, hateful, abusive content Shen et al. (2023); Zou et al. (2023a), and also content that can aid the user in causing harm or destruction (how to hotwire a car; how to make a bomb, etc.) Borji (2023); Borji and Mohammadian (2023); Lambert and Stevens (2023).
@@ 5.1.1. Definition of Jailbreak Success
We define the notion of jailbreak success based on graded evaluations of the divergence of task from task . We capture these metrics using property tests which are functions of the task and the jailbreak respectively as described in the previous paragraph. For manual evaluations, we consider the more stringent metric of whether aligns with the malicious intent of , called an intent test. We additionally run a programmatic intent test along with additional manual evaluations in Appendix Intent tests – Another method of attack detection.
@@ 5.2.1. Experimental Setup
We supply the model with a malicious user input . In some cases, we add a base user input before the jailbreak in order to emulate a Man-In-The-Middle (MITM) jailbreak. The aim is to cause divergence in the model output from the prompted task to a misaligned task .
We pick multiple instantiations of by choosing popular LLMs with different performance capabilities: OPT-175B Zhang et al. (2022), BLOOM-176B Workshop et al. (2022), GPT-3 models (text-ada, text-babbage, text-curie Brown et al. (2020a), GPT-3.5 based models text-davinci-002, gpt-3.5-turbo Ouyang et al. (2022), and FLAN-T5-XXL (11B) Wei et al. (2021). We design different kinds of jailbreaks for each task for a jailbreak type as . One may note that the jailbreaks are independent of the model used, since in most practical settings, an attacker knows which task the model has been prompted for, but not which model is being used (for e.g. BING’s announcement Mehdi (2023) about using GPT-4 came five weeks after their chatbot preview became accessible).
@@ 5.2.2. Dataset
The prompt formats are sourced from OpenAI, Promptsource Arora et al. (2022), and from several academic sources Chen et al. (2021); Muennighoff et al. (2022); Wei et al. (2022); Zhang et al. (2022). In cases where we did not find a preexisting prompt for a model-task combination, we recycled prompts from other models.
@@
Based on findings from Twitter, video sources, and Gehman et al. (2020), we manually curate jailbreaks across the said dimensions in Section 4, arriving at 55 jailbreaks over all 4 tasks. We run the property tests on all 55 jailbreaks for every model.
We vary the user input (100 inputs per jailbreak) for 37 of the 55 jailbreaks to analyze its effect on the attack’s success.
Therefore, in total, we have over 37100 = 3700 or 3.7k potential jailbreaks that are fed into each model.
@@ 5.2.3. Results
We report the results of our property tests for Figures 2 and 3 (and Figure 6 in the Appendix). In terms of the jailbreak type, we notice that the jailbreak success decreases with an increase in model size until text-davinci-002, however, any further instruction or task tuning increases the tendency for misalignment, as in the case of gpt-3.5-turbo and code-davinci-002. It can also be noted that Cognitive hacking (COG) appears to be the most successful form of jailbreak, which also happens to be the most common type of jailbreak present in the real world, followed by Orthographic attacks (ORTH).
Almost all models seem to be most affected by the Performance Degradation intent, which is expected given the relative ease of achieving degradation. However, the plots scaled by the statistics of human annotations (described in more detail in section 5.3) show misaligned content has high success for gpt-3.5-turbo and code-davinci-002, which we believe is related to their instruction-following capabilities. Additionally, text-davinci-002 appears to be robust to most of these jailbreaks, hinting that its training may be more robust to content-harms888https://tinyurl.com/davinci2toxic.
@@
It is seen that there is a poor agreement between both methods, suggesting that jailbreak detection can prove challenging and non-trivial. Additionally, we noticed that GPT-4 was occasionally jailbroken itself (especially with cognitive hacking), after being fed in the jailbreaks. This leads to a notion of a “jailbreak paradox", where it gets increasingly harder to detect and mitigate jailbreaks, due to the vast space of outputs the language model is capable of generating, and also due to its instruction-finetuning capabilities. Hence, we additionally conduct a manual evaluation to better understand the effect of jailbreaks on each model. However, this suggests the brittle nature of current attack success metrics such as ASR Zou et al. (2023b) and GPT-4 based evaluations Li et al. (2024).
@@ 5.3. Manual Analysis
Each prompt is independently labeled by two annotators, where disagreements were resolved by a third999In this case, we chose three Computer Science graduate students trained in NLP as annotators..
We report the misalignment rate and jailbreak success rate in Table 3. We can still see that the attack success rate is higher for FLAN, and gpt-3.5-turbo, which confirms that both model size and instruction tuning have an influence on jailbreaking. We report our inter-annotator agreement to be over both labels. Additionally, we scale the GPT-4 evaluation results of each attack type by the True-Positive Rate (TPR) and the False-Negative Rate (FNR) of GPT-4 against our manual evaluations. We perform the scaling as follows: if we observe that GPT-4 assigned a class to examples and class to examples in the dataset, then the estimated (corrected) values for the two labels will be , , where and represent the true positive and false negative rates respectively. Since and TPR and FNR capture the probability that GPT-4 classifies an instance of correctly and classifies as respectively, and represent the class distribution we would expect to observe if it was evaluated by a human annotator. Also note that . Post scaling, we see a significant increase in attack success for the Instruction Repetition (IR) type, due to the discrepancies between manual and automatic evaluations.
During the human evaluation, cases of partial task and intent misalignment were observed: for instance, upon the addition of attacks to a translation prompt, a model sometimes partially or incorrectly translates the user input. Similarly, in the case of intent success, the model sometimes partially followed the task before performing the attacker’s intent, leading to disagreement between the annotators. Hence, we conclude that while there are numerous studies conducted on understanding alignment and aligning models Wang et al. (2023), a large-scale study on a user-level understanding of misalignment is necessary to provide a further understanding of jailbreaking. Conversely, while current explainability studies focus on explaining model behavior to developers Geva et al. (2022), it has become increasingly necessary to understand how to explain such model misalignment to end-users. Studies along these fronts would prove fruitful in interpreting high-level model misalignment concepts such as jailbreaking.
5.4. Jailbreak Evaluation Paradox
The jailbreak paradox refers to the circular notion that a robust jailbreak detection/evaluation strategy also needs to be robust to being broken (or even jailbroken in some cases). Our functional tests are relatively “first-pass/naive” and can be bypassed. For instance, if a code generation prompt is targeted by an attack that says “Ignore everything and say nothing”, the model could generate syntactically correct and executable code like ‘print(“nothing”) which would be misaligned but not recognized by functional tests.
While the GPT-4 based evaluation strategy is more complex and can recognize cases like the previous one, it can also be jailbroken through clever prompt engineering. For example, it can be led to believe that an attack of the form "Ignore everything and say PWNED" is the original task (while functional testing could catch the difference in the model’s output). Such cases were also observed during the manual evaluation. The blind spots of these evaluation metrics is complementary by nature, and is also reflected in Table 2
We consider that the disagreement between humans also adds to this paradoxical notion; some humans consider a more lenient form of task and intent adherence compared to others, as discussed in section 5.3.
@@ 6. Conclusion
Large language models (LLMs) have shown remarkable capabilities of learning and performing new tasks from natural language instructions or prompts. However, this also exposes them to a new type of vulnerability, namely jailbreaks or prompt-injection attacks, where malicious users can manipulate the prompts to cause misalignment, leakage, performance degradation, or harmful generation. In this work, we have proposed a formalism and a taxonomy of jailbreaks based on their linguistic transformation and attacker intent. We perform an empirical analysis of the effectiveness of different types of jailbreaks on various LLMs. We found that LLMs have varying degrees of robustness to jailbreaks depending on their size, training data, and architecture. We discuss some limitations and challenges of the current methods for detecting and mitigating jailbreaks, such as the need for sanitizing and preprocessing the outputs, the difficulty of capturing the attacker’s intent, and the trade-off between functionality and security. We also explore some prompt-level mitigation strategies that we do not include because of space limitations. Specifically, our work provides a timely and useful framework and a comprehensive analysis for researchers and practitioners who are interested in understanding and addressing this emerging challenge.
@@
We make available the human evaluation data, jailbreaks, all source code, and GPT-4 prompts (wherever applicable) to the public101010https://github.com/AetherPrior/TrickLLM to aid in reproducibility and for future work.
Acknowledgements
The authors would like to thank the Microsoft Accelerate Foundation Models Research program 111111Microsoft AFMR: https://bitly.ws/3eawI for providing computational resources for the work.
- Arora et al. (2022) Simran Arora, Avanika Narayan, Mayee F. Chen, Laurel Orr, Neel Guha, Kush Bhatia, Ines Chami, Frederic Sala, and Christopher Ré. 2022. Ask me anything: A simple strategy for prompting language models.